# MIRI spectrophotometry of GN-z11: Detection and nature of an optical red continuum component

### v1

@@ 5 Summary and conclusions
This paper presents new imaging of GN-z11 with the MIRI F560W, F770W and F1000W filters, which extend the rest-frame UV-optical spectral coverage up to 0.86 m. In these MIRI images, GN-z11 shows a compact structure, compatible with previous results based on NIRCam imaging (Tacchella+23_GNz11). We have combined these new MIRI images with archival NIRSpec/Prism and MIRI/MRS spectroscopy, and NIRCam imaging to analyse the SED of GN-z11 covering the rest-frame 0.12 0.86 m. The complete spectral coverage of the NIRSpec and MRS spectroscopy allowed us to evaluate, and remove, the emission line contribution from the NIRCam and MIRI broad-band photometry, deriving the continuum-only SED of GN-z11. The main results drawn from this analysis can be summarised as follows:
@@
Classical colour-colour diagrams have been used to compare the observed colours with several SED templates including stellar populations of different ages, AGNs and high- LRDs. None of the templates can explain the colours of GN-z11 over the entire spectral range covered by our data. The rest-frame UV-optical colours (i.e. F200W F444W versus F444W F560W) of GN-z11 are compatible with the SEDs of unobscured young and intermediate age instantaneous bursts (including nebular emission), while not consistent with their optical and optical-red colours (i.e. F444W F560W, F560W F770W and F770W F1000W). The SEDs of high- LRDs and QSOs are not consistent with any of the UV-optical-red colours. The spectrum of Mrk 3, classified as a classical Seyfert 2, is compatible with the optical-red colours but fails to explain the observed UV-optical colours.
@@
The GN-z11 SED beyond 0.66 m cannot be entirely explained by the stellar and associated nebular emission alone. The observed flux in the F770W filter (rest-frame 0.66m) shows an excess of 36 3% above the predicted flux for a mixed stellar population of young (4 Myr) and intermediate (63 Myr) age stars. The flux marginally detected in the F1000W filter (rest-frame 0.86 m) represents an even larger excess, although with larger uncertainties (i.e., 91 28%).
@@
Both the continuum-only colours and the SED analysis covering the entire 0.12 0.86 m spectral range identify an optical-red continuum excess at wavelengths 0.66 0.86 m. This hints to a red source which appears undetected in the ultraviolet and represents a minor contribution to the emission in the optical range (i.e., 0.6 m). The origin of this excess is unclear. One scenario proposes that it can be produced by the emission from the dusty torus of a type 2 AGN characterised by a radius of 3 9 pc with a mass of (0.5 1.5) 105 . Alternatively, the excess could be related to hot dust emission or photoluminiscence of dust (Extended Red Emission) in the presence of the extreme radiation field (G/G0= 4.3105) produced by the compact and massive starburst in GN-z11, as already hypothesised in some low- low-metallicity starbursts. The presence of a standard type 1 AGN is not supported by the present data as an additional ad-hoc red source is required to explain the observed flux excess.
@@
In summary, flux excess has been identified in the MIRI/F770W (rest- 0.66 m) and also, although with lower significance, in MIRI/F1000W (rest- 0.86 m) images of GN-z11. This excess is likely due to the dust emission associated to an AGN dusty torus or to dust in the environment of an extremely compact and young starburst. However, the current wavelength coverage is still limited and the true nature of the heating source is still unclear. Deep imaging with MIRI/F1000W and redder filters are required to confirm the 0.86 m detection and extend the SED coverage beyond H, characterizing its near-IR spectral shape up to at least 1.1 m. Imaging with deep integrations using MIRI F1000W and F1280W filters will be executed with JWST in the near future (PID 7605). This will help to discriminate between the two proposed scenarios. In addition, deeper than existing MIRI/MRS spectroscopy would also be required to detect or rule out the presence of a weak broad H associated with a 106 type 1 AGN, or to identify the existence of dusty outflows.
@@ Acknowledgements.
A.C.G. acknowledges support by JWST contract B0215/JWST-GO-02926. L.C. and J.A-M. acknowledge support by grants PIB2021-127718NB-I00 and PID2024-158856NA-I00, A.A-H. by grant PID2021-124665NB-I00, and P.G.P-G by grant PID2022-139567NB-I00 funded by the Spanish Ministry of Science and Innovation and the State Agency of Research MCIN/AEI/10.13039/501100011033 and ERDF ”A way of making Europe”. M.A. acknowledges financial support from Comunidad de Madrid under Atracción de Talento grant 2020-T2/TIC-19971. D.L. was supported by research grants (VIL16599, VIL54489) from VILLUM FONDEN. A.B. and G.O. acknowledge support from the Swedish National Space Administration (SNSA). AJB acknowledges funding from the ”FirstGalaxies” Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 789056). L.A.B. acknowledges support from the Dutch Research Council (NWO) under grant VI.Veni.242.055 (https://doi.org/10.61686/LAJVP77714) and the ERC Consolidator grant 101088676 (”VOYAJ”). This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST; and from the European JWST archive (eJWST) operated by the ESDC.Appendix A NIRCam-NIRSpec-MIRI normalization

### v2

@@ 5 Summary and conclusions
This paper presents new imaging of GN-z11 with the MIRI F560W, F770W, and F1000W filters, which extend the rest-frame UV-optical coverage up to 0.86 m. In these MIRI images, GN-z11 shows a compact structure, compatible with previous results based on NIRCam imaging (Tacchella et al., 2023a). We have combined these new MIRI images with archival NIRSpec/Prism and MIRI/MRS spectroscopy, and NIRCam imaging, to analyse the SED of GN-z11 covering the rest-frame 0.12 0.86 m. In addition, the completeness of the spectroscopic NIRSpec and MRS dataset allowed us to remove the emission line contribution from the NIRCam and MIRI broad-band photometry, deriving the SED of GN-z11 with line-free, continuum-only fluxes. The main results drawn from this analysis can be summarised as follows:
@@
Classical colour-colour diagrams have been used to compare the observed colours with several SED templates including stellar populations of different ages, AGNs, and high- LRDs. None of the templates can explain the colours of GN-z11 over the entire spectral range covered by our data. The rest-frame UV-optical colours (i.e. F200W F444W versus F444W F560W) of GN-z11 are compatible with the SEDs of unobscured young and intermediate age instantaneous bursts (including nebular emission), while not consistent with their optical and optical-red colours (i.e. F444W F560W, F560W F770W and F770W F1000W). The SEDs of high- LRDs and QSOs are not consistent with any of the UV-optical-red colours. The spectrum of Mrk 3, classified as a classical Seyfert 2, is compatible with the optical-red colours but fails to explain the observed UV-optical colours.
@@
The GN-z11 SED beyond 0.66 m cannot be entirely explained by the stellar and associated nebular emission alone. The observed flux in the F770W filter (rest-frame 0.66m) shows an excess of 36 3% above the predicted flux for a mixed stellar population of young (4 Myr) and intermediate-age (63 Myr) stars. The flux marginally detected in the F1000W filter (rest-frame 0.86 m) represents an even larger excess, although one with larger uncertainties (i.e. 91 28%).
@@
Both the continuum-only colours and the SED analysis covering the entire 0.12 0.86 m spectral range identify a optical-red continuum excess at wavelengths 0.66 0.86 m. This hints at a red source that appears undetected in the ultraviolet and represents a minor contribution to the emission in the optical range (i.e. 0.6 m). Although the origin of this excess is unclear, we propose that it could be produced by the emission from the dusty torus of a type 2 AGN characterised by a radius of 3 9 pc with a mass of (0.5 1.5) 105 . Alternatively, the excess could be related to hot dust emission or photoluminiscence of dust (Extended Red Emission) in the presence of the extreme radiation field (G/G0= 4.3105) produced by the compact and massive starburst in GN-z11, as has already been hypothesised for some low- low-metallicity starbursts. The presence of a standard type 1 AGN is not supported by the present data as an additional ad hoc red source is required to explain the observed flux excess.
@@
In summary, flux excess has been identified in the MIRI/F770W (rest- 0.66 m) and also, although with lower significance, in MIRI/F1000W (rest- 0.86 m) images of GN-z11. This excess is likely due to the dust emission associated with an AGN dusty torus or with dust in the environment of an extremely compact and young starburst. However, the current wavelength coverage is still limited and the true nature of the heating source is still unclear. Deep imaging with MIRI/F1000W and redder filters are required to confirm the 0.86 m detection and extend the SED coverage beyond H, characterising its near-IR spectral shape up to at least 1.1 m. Imaging with deep integrations using MIRI F1000W and F1280W filters will be executed with JWST in the near future (PID 7605). This will help to discriminate between the two proposed scenarios. In addition, deeper MIRI/MRS spectroscopy than is in existence would also be required to detect or rule out the presence of a weak broad H associated with a 106 type 1 AGN, or identify the existence of dusty outflows.
@@ Acknowledgements.
A.C.G. acknowledges support by JWST contract B0215/JWST-GO-02926. L.C. and J.A-M. acknowledge support by grants PIB2021-127718NB-I00 and PID2024-158856NA-I00, A.A-H. by grant PID2021-124665NB-I00, and P.G.P-G by grant PID2022-139567NB-I00 funded by the Spanish Ministry of Science and Innovation and the State Agency of Research MCIN/AEI/10.13039/501100011033 and ERDF ”A way of making Europe”. M.A. acknowledges financial support from Comunidad de Madrid under Atracción de Talento grant 2020-T2/TIC-19971. D.L. was supported by research grants (VIL16599, VIL54489) from VILLUM FONDEN. A.B. and G.O. acknowledge support from the Swedish National Space Administration (SNSA). AJB acknowledges funding from the ”FirstGalaxies” Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 789056). L.A.B. acknowledges support from the Dutch Research Council (NWO) under grant VI.Veni.242.055 (https://doi.org/10.61686/LAJVP77714) and the ERC Consolidator grant 101088676 (”VOYAJ”). This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST; and from the European JWST archive (eJWST) operated by the ESDC.References
- Super star clusters in Haro 11: properties of a very young starburst and evidence for a near-infrared flux excess. MNRAS 407 (2), pp. 870–890. External Links: Document, 1005.1658, ADS entry Cited by: §4.4.2.
- On the Origin of the Red Excess in Very Young Super Star Clusters: The Case of SBS 0335-052E. ApJ 725 (2), pp. 1620–1628. External Links: Document, 1010.3703, ADS entry Cited by: §4.4.2, §4.4.2.
- COSMOS-Web: The Overabundance and Physical Nature of “Little Red Dots”—Implications for Early Galaxy and SMBH Assembly. ApJ 991 (1), pp. 37. External Links: Document, 2406.10341, ADS entry Cited by: Figure 2, Figure 3, §4.1.
- Insight into the starburst nature of Galaxy GN-z11 with JWST MIRI spectroscopy. A&A 695, pp. A250. External Links: Document, 2412.12826, ADS entry Cited by: §1, §1, §2.2, §3.2, §3.2, §3.3, §4.2, §4.3, §4.4.1, §4.4.1, §4.4.2, §4.4.2, footnote 5.
- MIRI/JWST observations reveal an extremely obscured starburst in the z = 6.9 system SPT0311-58. A&A 671, pp. A105. External Links: Document, 2301.02313, ADS entry Cited by: §2.1.
- JWST MIRI flight performance: The Medium-Resolution Spectrometer. A&A 675, pp. A111. External Links: Document, 2303.13469, ADS entry Cited by: §2.2.
- New Clues on the Extended He II Ionization in IZw18 from GTC/MEGARA and JWST/MIRI. ApJ 987 (2), pp. L36. External Links: Document, ADS entry Cited by: §4.4.2.
- Is GN-z11 powered by a super-Eddington massive black hole?. A&A 686, pp. A141. External Links: Document, 2401.13733, ADS entry Cited by: §1.
- CIGALE: a python Code Investigating GALaxy Emission. A&A 622, pp. A103. External Links: Document, 1811.03094, ADS entry Cited by: §3.3.
- A mature quasar at cosmic dawn revealed by JWST rest-frame infrared spectroscopy. Nature Astronomy 8, pp. 1054–1065. External Links: Document, 2307.14414, ADS entry Cited by: Figure 2, Figure 3, §4.1, §4.3, §4.4.1.
- A close look at the black hole masses and hot dusty toruses of the first quasars with MIRI-MRS. sub. ApJ, pp. arXiv:2511.02902. External Links: Document, 2511.02902, ADS entry Cited by: §3.3, Figure 2, Figure 3, §4.1.
- The Mid-Infrared Instrument for the James Webb Space Telescope, III: MIRIM, The MIRI Imager. PASP 127 (953), pp. 612. External Links: Document, 1508.02488, ADS entry Cited by: §2.1.
- z ~7 Galaxy Candidates from NICMOS Observations Over the HDF-South and the CDF-South and HDF-North Goods Fields. ApJ 725 (2), pp. 1587–1599. External Links: Document, 1003.1706, ADS entry Cited by: §1.
- Stellar population synthesis at the resolution of 2003. MNRAS 344 (4), pp. 1000–1028. External Links: Document, astro-ph/0309134, ADS entry Cited by: §3.3.
- JADES NIRSpec initial data release for the Hubble Ultra Deep Field: Redshifts and line fluxes of distant galaxies from the deepest JWST Cycle 1 NIRSpec multi-object spectroscopy. A&A 690, pp. A288. External Links: Document, 2306.02467, ADS entry Cited by: §2.2.
- JADES NIRSpec Spectroscopy of GN-z11: Lyman- emission and possible enhanced nitrogen abundance in a z = 10.60 luminous galaxy. A&A 677, pp. A88. External Links: Document, 2302.07256, ADS entry Cited by: Appendix A, §1, §1, §2.2, §3.2, §3.2, §3.3, Figure 5, §4.3, §4.4.2, §4.4.2.
- Star formation and dust attenuation properties in galaxies from a statistical ultraviolet-to-far-infrared analysis. MNRAS 360 (4), pp. 1413–1425. External Links: Document, astro-ph/0504434, ADS entry Cited by: §3.3.
- The Brightest Young Star Clusters in NGC 5253.. ApJ 811 (2), pp. 75. External Links: Document, 1508.04476, ADS entry Cited by: §4.4.2.
- The Dust Content and Opacity of Actively Star-forming Galaxies. ApJ 533 (2), pp. 682–695. External Links: Document, astro-ph/9911459, ADS entry Cited by: §3.3, §3.3, §3.3.
- JADES: Probing interstellar medium conditions at z 5.5-9.5 with ultra-deep JWST/NIRSpec spectroscopy. A&A 677, pp. A115. External Links: Document, 2302.04298, ADS entry Cited by: §4.4.2.
- Spectroscopic confirmation of two luminous galaxies at a redshift of 14. Nature 633 (8029), pp. 318–322. External Links: Document, 2405.18485, ADS entry Cited by: Appendix A, §1.
- Galactic Stellar and Substellar Initial Mass Function. PASP 115 (809), pp. 763–795. External Links: Document, astro-ph/0304382, ADS entry Cited by: §1, §3.3, §3.3.
- N-enhancement in GN-z11: First evidence for supermassive stars nucleosynthesis in proto-globular clusters-like conditions at high redshift?. A&A 673, pp. L7. External Links: Document, 2303.07955, ADS entry Cited by: §1.
- A Simple Model for the Absorption of Starlight by Dust in Galaxies. ApJ 539 (2), pp. 718–731. External Links: Document, astro-ph/0003128, ADS entry Cited by: §3.3.
- RIOJA. Dusty outflows and density-complex ISM in the N-enhanced lensed galaxy RXCJ2248-ID at z=6.1. sub. A&A, pp. arXiv:2511.14658. External Links: Document, 2511.14658, ADS entry Cited by: §4.4.2.
- The mass-metallicity and the fundamental metallicity relation revisited on a fully Te-based abundance scale for galaxies. MNRAS 491 (1), pp. 944–964. External Links: Document, 1910.00597, ADS entry Cited by: §3.2.
- Spectroscopic confirmation of four metal-poor galaxies at z = 10.3-13.2. Nature Astronomy 7, pp. 622–632. External Links: Document, 2212.04568, ADS entry Cited by: §1.
- JADES Data Release 3: NIRSpec/Microshutter Assembly Spectroscopy for 4000 Galaxies in the GOODS Fields. ApJS 277 (1), pp. 4. External Links: Document, 2404.06531, ADS entry Cited by: §2.2.
- A remarkable ruby: Absorption in dense gas, rather than evolved stars, drives the extreme Balmer break of a little red dot at z = 3.5. A&A 701, pp. A168. External Links: Document, 2503.16600, ADS entry Cited by: Figure 2, Figure 3, §4.1.
- Efficient formation of massive galaxies at cosmic dawn by feedback-free starbursts. MNRAS 523 (3), pp. 3201–3218. External Links: Document, 2303.04827, ADS entry Cited by: §4.4.2.
- The JADES Origins Field: A New JWST Deep Field in the JADES Second NIRCam Data Release. Accepted to ApJ Supplement, pp. arXiv:2310.12340. External Links: Document, 2310.12340, ADS entry Cited by: §1, §2.2.
- The 2017 Release Cloudy. Rev. Mexicana Astron. Astrofis. 53, pp. 385–438. External Links: Document, 1705.10877, ADS entry Cited by: §3.3, §4.1.
- On the stunning abundance of super-early, luminous galaxies revealed by JWST. MNRAS 522 (3), pp. 3986–3991. External Links: Document, 2208.00720, ADS entry Cited by: §4.4.2.
- Dusty-wind-clear JWST Super-early Galaxies. ApJ 943 (2), pp. L27. External Links: Document, 2211.08937, ADS entry Cited by: §4.4.2.
- Revisiting the infrared spectra of active galactic nuclei with a new torus emission model. MNRAS 366 (3), pp. 767–786. External Links: Document, astro-ph/0511428, ADS entry Cited by: §3.3, Figure 5.
- The NOEMA observations of GN-z11: constraining the neutral interstellar medium and dust formation in the heart of cosmic reionization at z = 10.6. MNRAS 530 (1), pp. 340–346. External Links: Document, 2309.02493, ADS entry Cited by: §1.
- Torus model properties of an ultra-hard X-ray selected sample of Seyfert galaxies. MNRAS 486 (4), pp. 4917–4935. External Links: Document, 1904.03694, ADS entry Cited by: §4.4.1, §4.4.1.
- Outshining in the spatially resolved analysis of a strongly lensed galaxy at z = 6.072 with JWST NIRCam. A&A 686, pp. A63. External Links: Document, 2402.17875, ADS entry Cited by: §4.2.
- Spatially Resolved Properties of Galaxies at 5 ¡ z ¡ 9 in the SMACS 0723 JWST ERO Field. ApJ 948 (2), pp. 126. External Links: Document, 2212.08670, ADS entry Cited by: §4.2.
- A Near-Infrared Spectral Template for Quasars. ApJ 640 (2), pp. 579–591. External Links: Document, astro-ph/0511640, ADS entry Cited by: §3.3, Figure 2, Figure 3, §4.1, §4.4.1.
- Estimating Black Hole Masses in Active Galaxies Using the H Emission Line. ApJ 630 (1), pp. 122–129. External Links: Document, astro-ph/0508335, ADS entry Cited by: §1.
- The nature of the torus in the heavily obscured AGN Markarian 3: an X-ray study. MNRAS 460 (2), pp. 1954–1969. External Links: Document, 1605.02467, ADS entry Cited by: §4.1.
- The 23.01 Release of Cloudy. Research Notes of the American Astronomical Society 7 (11), pp. 246. External Links: Document, 2311.10163, ADS entry Cited by: §3.3.
- JADES NIRSpec spectroscopy of GN-z11: evidence for Wolf–Rayet contribution to stellar populations at 430 Myr after big bang?. MNRAS 543 (4), pp. 3172–3195. External Links: Document, 2504.12584, ADS entry Cited by: §4.4.2.
- The interstellar radiation density between 912 A and 2400 A. Bull. Astron. Inst. Netherlands 19, pp. 421. External Links: ADS entry Cited by: §4.4.2.
- Imaging of I Zw 18 by JWST. I. Detecting Dusty Stellar Populations. AJ 168 (1), pp. 23. External Links: Document, 2403.06980, ADS entry Cited by: §4.4.2.
- Detection of Nuclear X-Ray Sources in Nearby Galaxies with Chandra. ApJ 549 (1), pp. L51–L54. External Links: Document, astro-ph/0102504, ADS entry Cited by: §4.4.1.
- Detection of Nuclear X-Ray Sources in Nearby Galaxies with Chandra. ApJ 549 (1), pp. L51–L54. External Links: Document, astro-ph/0102504, ADS entry Cited by: §1.
- JADES – the small blue bump in GN-z11: insights into the nuclear region of a galaxy at z = 10.6. MNRAS 541 (3), pp. 2134–2161. External Links: Document, 2405.05772, ADS entry Cited by: §1.
- Evidence for GN-z11 as a luminous galaxy at redshift 10.957. Nature Astronomy 5, pp. 256–261. External Links: Document, 2012.06936, ADS entry Cited by: §4.1.
- The extended He II 4686 emission in the extremely metal-poor galaxy SBS 0335 - 052E seen with MUSE. MNRAS 480 (1), pp. 1081–1095. External Links: Document, 1807.09307, ADS entry Cited by: §4.4.2.
- Using Strong Lines to Estimate Abundances in Extragalactic H II Regions and Starburst Galaxies. ApJS 142 (1), pp. 35–52. External Links: Document, astro-ph/0206495, ADS entry Cited by: §3.2.
- Resolving Emission from Small Dust Grains in the Blue Compact Dwarf II Zw 40 with JWST. ApJ 991 (2), pp. L56. External Links: Document, 2509.04662, ADS entry Cited by: §4.4.2.
- From Dust to Nanodust: Resolving Circumstellar Dust from the Colliding-wind Binary Wolf-Rayet 140. ApJ 951 (2), pp. 89. External Links: Document, 2305.14557, ADS entry Cited by: §4.4.2.
- The VIMOS Ultra-Deep Survey: evidence for AGN feedback in galaxies with CIII]-1908 Å emission 10.8 to 12.5 Gyr ago. A&A 625, pp. A51. External Links: Document, 1710.10715, ADS entry Cited by: §1.
- Physical Properties of II Zw 40’s Super Star Cluster and Nebula: New Insights and Puzzles from UV Spectroscopy. ApJ 865 (1), pp. 55. External Links: Document, 1808.04332, ADS entry Cited by: §4.4.2.
- Feedback-free starbursts at cosmic dawn: Observable predictions for JWST. A&A 690, pp. A108. External Links: Document, 2311.14662, ADS entry Cited by: §4.4.2.
- High-precision Astrometry and Photometry with the JWST/MIRI Imager. PASP 136 (3), pp. 034502. External Links: Document, 2311.12145, ADS entry Cited by: Appendix A, §3.1.
- A small and vigorous black hole in the early Universe. Nature 627 (8002), pp. 59–63. External Links: Document, 2305.12492, ADS entry Cited by: §1, §1, §4.3, §4.4.1.
- Dust Formation in Massive WR+O Binaries: Recent Results. In Massive Stars in Interactive Binaries, N. St. -Louis and A. F. J. Moffat (Eds.), Astronomical Society of the Pacific Conference Series, Vol. 367, pp. 213. External Links: Document, astro-ph/0610531, ADS entry Cited by: §4.4.2.
- Local supermassive black holes, relics of active galactic nuclei and the X-ray background. MNRAS 351 (1), pp. 169–185. External Links: Document, astro-ph/0311619, ADS entry Cited by: §4.4.1.
- Extremely UV-bright starbursts at the end of cosmic reionization. arXiv e-prints, pp. arXiv:2510.12411. External Links: Document, 2510.12411, ADS entry Cited by: §4.4.2.
- Narrow-line AGN selection in CEERS: Spectroscopic selection, physical properties, and X-ray and radio analysis. A&A 700, pp. A12. External Links: Document, 2408.15615, ADS entry Cited by: §4.1.
- New AGN diagnostic diagrams based on the [OIII]4363 auroral line. A&A 691, pp. A345. External Links: Document, 2404.10811, ADS entry Cited by: §4.1.
- Spectropolarimetry of High-Polarization Seyfert 2 Galaxies and Unified Seyfert Theories. ApJ 355, pp. 456. External Links: Document, ADS entry Cited by: §4.1.
- Exploring the Mysterious High-ionization Source Powering [Ne V] in High-z Analog SBS0335-052 E with JWST/MIRI. ApJ 985 (2), pp. 253. External Links: Document, 2502.07662, ADS entry Cited by: §4.4.2.
- A “Black Hole Star” Reveals the Remarkable Gas-Enshrouded Hearts of the Little Red Dots. arXiv e-prints, pp. arXiv:2503.16596. External Links: Document, 2503.16596, ADS entry Cited by: §1, §4.1.
- AGN Dusty Tori. I. Handling of Clumpy Media. ApJ 685 (1), pp. 147–159. External Links: Document, 0806.0511, ADS entry Cited by: §4.4.1.
- AGN Dusty Tori. II. Observational Implications of Clumpiness. ApJ 685 (1), pp. 160–180. External Links: Document, 0806.0512, ADS entry Cited by: §4.4.1.
- Analysis of galaxy spectral energy distributions from far-UV to far-IR with CIGALE: studying a SINGS test sample. A&A 507 (3), pp. 1793–1813. External Links: Document, 0909.5439, ADS entry Cited by: §3.3.
- A Remarkably Luminous Galaxy at z=11.1 Measured with Hubble Space Telescope Grism Spectroscopy. ApJ 819 (2), pp. 129. External Links: Document, 1603.00461, ADS entry Cited by: §1.
- MIRI Deep Imaging Survey (MIDIS) of the Hubble Ultra Deep Field: Survey description and early results for the galaxy population detected at 5.6 m. A&A 696, pp. A57. External Links: Document, 2411.19686, ADS entry Cited by: §2.1, §3.1.
- Stellar populations in local star-forming galaxies - I. Data and modelling procedure. MNRAS 338 (2), pp. 508–524. External Links: Document, astro-ph/0209396, ADS entry Cited by: §3.3.
- What Is the Nature of Little Red Dots and what Is Not, MIRI SMILES Edition. ApJ 968 (1), pp. 4. External Links: Document, 2401.08782, ADS entry Cited by: §2.1.
- What Is the Nature of Little Red Dots and what Is Not, MIRI SMILES Edition. ApJ 968 (1), pp. 4. External Links: Document, 2401.08782, ADS entry Cited by: Figure 2, Figure 3, §4.1.
- The Stellar Mass Assembly of Galaxies from z = 0 to z = 4: Analysis of a Sample Selected in the Rest-Frame Near-Infrared with Spitzer. ApJ 675 (1), pp. 234–261. External Links: Document, 0709.1354, ADS entry Cited by: §3.3.
- Planck 2018 results. VI. Cosmological parameters. A&A 641, pp. A6. External Links: Document, 1807.06209, ADS entry Cited by: §1.
- Spectral Energy Distributions of Hard X-Ray Selected Active Galactic Nuclei in the XMM-Newton Medium Deep Survey. ApJ 663 (1), pp. 81–102. External Links: Document, astro-ph/0703255, ADS entry Cited by: §3.3.
- Chandra and Spitzer Unveil Heavily Obscured Quasars in the Chandra/SWIRE Survey. ApJ 642 (2), pp. 673–693. External Links: Document, astro-ph/0602228, ADS entry Cited by: §3.3.
- a New View of the Super Star Clusters in the Low-Metallicity Galaxy SBS 0335-052. AJ 136 (4), pp. 1415–1426. External Links: Document, 0807.1098, ADS entry Cited by: §4.4.2, §4.4.2.
- Identification and properties of intense star-forming galaxies at redshifts z ¿ 10. Nature Astronomy 7, pp. 611–621. External Links: Document, 2212.04480, ADS entry Cited by: §1.
- GN-z11: The environment of an active galactic nucleus at z = 10.603. New insights into the most distant Ly detection. A&A 687, pp. A283. External Links: Document, 2306.09142, ADS entry Cited by: §1.
- An X-Shooter composite of bright 1 ¡ z ¡ 2 quasars from UV to infrared. A&A 585, pp. A87. External Links: Document, 1510.08058, ADS entry Cited by: §3.3, Figure 2, Figure 3, §4.1, §4.4.1.
- GN-z11 in Context: Possible Signatures of Globular Cluster Precursors at Redshift 10. ApJ 966 (1), pp. 92. External Links: Document, ADS entry Cited by: §1.
- CLusters in the UV as EngineS (CLUES). I. Survey Presentation and FUV Spectral Analysis of the Stellar Light. AJ 164 (5), pp. 208. External Links: Document, 2209.09914, ADS entry Cited by: §4.4.2.
- An Atlas of Hubble Space Telescope STIS Spectra of Seyfert Galaxies. ApJS 166 (2), pp. 498–504. External Links: Document, astro-ph/0605673, ADS entry Cited by: Figure 2, Figure 3, §4.1.
- 3D radiative transfer modelling of the dusty tori around active galactic nuclei as a clumpy two-phase medium. MNRAS 420 (4), pp. 2756–2772. External Links: Document, 1109.1286, ADS entry Cited by: §4.3.
- The dust covering factor in active galactic nuclei. MNRAS 458 (3), pp. 2288–2302. External Links: Document, 1602.06954, ADS entry Cited by: §4.3.
- Re-evaluating old stellar populations. MNRAS 479 (1), pp. 75–93. External Links: Document, 1805.08784, ADS entry Cited by: §3.3.
- JADES Imaging of GN-z11: Revealing the Morphology and Environment of a Luminous Galaxy 430 Myr after the Big Bang. ApJ 952 (1), pp. 74. External Links: Document, 2302.07234, ADS entry Cited by: Appendix A, Appendix A, §1, §1, §4.1, §4.2, §4.3, §4.3, §4.4.2, §5.
- JWST NIRCam + NIRSpec: interstellar medium and stellar populations of young galaxies with rising star formation and evolving gas reservoirs. MNRAS 522 (4), pp. 6236–6249. External Links: Document, 2208.03281, ADS entry Cited by: §4.2.
- CAPERS-LRD-z9: A Gas-enshrouded Little Red Dot Hosting a Broad-line Active Galactic Nucleus at z = 9.288. ApJ 989 (1), pp. L7. External Links: Document, 2505.04609, ADS entry Cited by: §4.1.
- Multiline Assessment of Narrow-line Regions in z 3 Radio Galaxies. ApJ 929 (1), pp. 51. External Links: Document, 2203.03390, ADS entry Cited by: §1.
- The Spitzer View of Low-Metallicity Star Formation. II. Mrk 996, a Blue Compact Dwarf Galaxy with an Extremely Dense Nucleus. ApJ 689 (2), pp. 897–912. External Links: Document, 0809.3082, ADS entry Cited by: §4.4.2.
- Metal-poor star formation at z ¿ 6 with JWST: new insight into hard radiation fields and nitrogen enrichment on 20 pc scales. MNRAS 529 (4), pp. 3301–3322. External Links: Document, 2401.08764, ADS entry Cited by: §4.4.2.
- Dust and super star clusters in NGC 5253. A&A 415, pp. 509–520. External Links: Document, astro-ph/0312364, ADS entry Cited by: §4.4.2.
- Extended Red Emission: Photoluminescence by Interstellar Nanoparticles. In Astrophysics of Dust, A. N. Witt, G. C. Clayton, and B. T. Draine (Eds.), Astronomical Society of the Pacific Conference Series, Vol. 309, pp. 115. External Links: Document, astro-ph/0309674, ADS entry Cited by: §4.4.2.
- Stars and gas in the most metal-poor galaxies - I. COS and MUSE observations of SBS 0335-052E. MNRAS 500 (3), pp. 2908–2927. External Links: Document, 2010.13963, ADS entry Cited by: §4.4.2.
- X-CIGALE: Fitting AGN/galaxy SEDs from X-ray to infrared. MNRAS 491 (1), pp. 740–757. External Links: Document, 2001.08263, ADS entry Cited by: §4.4.1.
- Blue monsters. Why are JWST super-early, massive galaxies so blue?. MNRAS 520 (2), pp. 2445–2450. External Links: Document, 2209.06840, ADS entry Cited by: §4.4.2.
Appendix A NIRCam-NIRSpec-MIRI normalisation